equation of the line in slope intercept form with a slope 2 and passes through (1,5)
Answer by josgarithmetic(39617)

Slope Intercept Form is y=mx+b.
m is slope, b is y intercept.

b=y-mx
b=5-2*1
b=3

Equation is therefore y=2x+3